Parth (parthkohli):

Just like indefinite integration is the inverse of differentiation, what is definite integration the inverse of?

OpenStudy (shubhamsrg):

definite integration is just the sub part of indefinite integration i'd guess..

Parth (parthkohli):

So there isn't an inverse of definite integration?

OpenStudy (shubhamsrg):

not in my knowledge,,

OpenStudy (shubhamsrg):

definite integration yields just another real no.

hartnn (hartnn):

i think it would be inverse of ' derivative at an end-point(upper limit)-derivative at another end-point(lower limit)' ....

Parth (parthkohli):

Haha, that really does make sense :)

hartnn (hartnn):

\[\huge{\frac{d}{dx}\int\limits_{a}^{b}}f(x)dx=f(b)-f(a)\] is this true ?

OpenStudy (shubhamsrg):

f(b) * b' - f(a) * a' it should be.. if a and b are constants..derivative is 0,ofcorse..